Silver metal reacts with sulfur to form silver sulfide according to the following balanced equation: (5 points)2Ag(𝑠) + S(𝑠) → 𝐴𝑔2S(𝑠)What is the limiting reactant when 50.0 g Ag is reacted with 10.0 g S?

1
Expert's answer
2022-04-12T14:01:03-0400

Solution:

The molar mass of silver (Ag) is 107.868 g/mol

The molar mass of sulfur (S) is 32.065 g/mol


Calculate the moles of each reactant:

Moles of Ag = (50.0 g Ag) × (1 mol Ag / 107.868 g Ag) = 0.464 mol Ag

Moles of S = (10.0 g S) × (1 mol S / 32.065 g S) = 0.312 mol S


Balanced chemical equation:

2Ag(s) + S(s) → Ag2S(s)

According to stoichiometry:

2 mol of Ag react with 1 mol of S

Thus, 0.464 mol of Ag react with:

(0.464 mol Ag) × (1 mol S / 2 mol Ag) = 0.232 mol S

However, initially there is 0.312 mol of S (according to the task)

Therefore, Ag acts as limiting reactant and S is excess reactant


Answer: Silver (Ag) is the limiting reactant